OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2?

OpenStudy (anonymous):

@koook do you think you can help me?

OpenStudy (anonymous):

@blurbendy

OpenStudy (anonymous):

point slope form (y-y1) = m(x-x1) where (x1,y1) are the given points and m = slope

OpenStudy (anonymous):

y+8=-2(x-6) is this correct?

OpenStudy (anonymous):

correct. :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

Do you think, you can help me with a few more?

OpenStudy (anonymous):

Sure.